# The Order Effect: Investigating Prompt Sensitivity in Closed-Source LLMs

### v1

@@ Abstract
As large language models (LLMs) become integral to diverse applications, ensuring their reliability under varying input conditions is crucial. One key issue affecting this reliability is order sensitivity, wherein slight variations in input arrangement can lead to inconsistent or biased outputs. Although recent advances have reduced this sensitivity, the problem remains unresolved. This paper investigates the extent of order sensitivity in closed-source LLMs by conducting experiments across multiple tasks, including paraphrasing, relevance judgment, and multiple-choice questions. Our results show that input order significantly affects performance across tasks, with shuffled inputs leading to measurable declines in output accuracy. Few-shot prompting demonstrates mixed effectiveness and offers partial mitigation, however, fails to fully resolve the problem. These findings highlight persistent risks, particularly in high-stakes applications, and point to the need for more robust LLMs or improved input-handling techniques in future development.
@@ 1 Introduction
One necessary requirement for reasoning is order independence. A model should provide the same consistent response to a query regardless of the order of its content. Historically, LLMs have struggled with this issue. Swapping subsequences within semantically identical inputs often leads to significant changes in output, a problem that worsens as inputs grow in size and complexity (He et al., 2024). Today, improvements in LLMs promise more accurate responses that mitigate order dependency. However, it still remains unclear whether these improvements are sufficient to reduce order dependency when these models are used in the wild.
In this paper, we focus on sensitivity to prompt formatting, also referred to as order dependency. This problem has been previously explored in the context of multiple-choice questions (Pezeshkpour & Hruschka, 2023; Zheng et al., 2024), laying the foundation for our research. Building on this foundation, we aim is to provide a fresh perspective and expand the analysis with additional data points to investigate the problem more thoroughly. Although the issue may seem trivial or some may question the need for further study, we demonstrate that it persists, continues to cause problems, and warrants ongoing investigation. We hope to uncover patterns or root causes that can help mitigate order dependency in the future.
To demonstrate the severity of the problem, we conducted a simple experiment in which we submitted the following prompt twice to GPT-4o mini:
@@
To investigate this problem, we experimented with GPT-4o and GPT-4o mini and measured their performance on various order-sensitive tasks. We chose to work with closed-source LLMs, specifically the GPT models, because they are widely used by non-technical users. As LLMs become increasingly integrated into society, many people who rely on them may lack the expertise to interact with them and assess their accuracy. Consequently, such interactions could contribute to the spread of inaccurate data across the internet, public discourse, and even scientific publications. The more we analyze and refine our understanding of these models, the better we can guide their use.
Although similar issues have been reported in open-source models, sometimes even more severe than in their closed-source counterparts, we are less concerned about their uses and thus exclude them from the scope of this paper. Open-source models are primarily used and developed by technical users who engage with them in a more controlled and reliable way. In contrast, closed-source models are accessible to a broader audience, which makes it crucial to understand their potential inconsistencies. We hope this paper contributes to the ongoing discussion regarding the reliability of LLMs as decision-making tools in complex and diverse domains.
@@ 3 Methodology and Experimental Design
To study the impact of order, we conduct five experiments, each consisting of four sub-experiments. In the first sub-experiment, we evaluate the models’ performance in a zero-shot setting using the original order of questions and corresponding choices, where the LLM’s task is to select the correct option. In the second sub-experiment, we again assess the models’ zero-shot performance, but this time the entries are presented in a randomized order. For example, a question with the original order of choices, such as “a” and “b”, will be presented in its new form with the choices reversed, namely “b” and “a” (see the following sections for prompt examples).
@@
Finally, in the fourth sub-experiment, we assess the models’ performance in a few-shot setting where the order of entries is randomized. This reordering does not follow any specific pattern. Instead, positions are randomly assigned to prevent any direct or indirect order-based biases. Together, these four setups allow us to compare model performance in zero-shot and few-shot settings and evaluate the impact of input order on their outputs.
@@ 3.1 Experiment 1: MRPC
Where sentence_1 and sentence_2 are variables which are replaced with the real sentences shared above. This prompt is used for the zero-shot setup with the original order. For the zero-shot setup with the shuffled order, the only change we make to the prompt is swapping the order of the sentences. Specifically, the part of the prompt needs to be modified is shown below:
@@
3.2 Experiment 2: WebGPT
The second experiment focuses on comparison consistency. We evaluate the models’ ability to compare two answers to a given question and determine which answer a human judge would prefer, regardless of the order in which the answers are presented. We use OpenAI’s WebGPT dataset (Nakano et al., 2021), [^2] https://huggingface.co/datasets/openai/webgpt_comparisons which provides pairwise comparisons derived from a reward model trained on human feedback to reflect real-world preferences for long-form question answering. To capture the models’ behavior, we switch the order of the answers and observe whether they consistently select the same preferred answer. We use the $1958$ examples from the test set. Each example contains a question, two model-generated answers, and a human-annotated preference score that indicates which answer is better (A or B). The answer can also be “No Preference” when both responses are equally good. To ensure that the prompt in the few-shot setting is informative, we include one example from each case (A is better than B, B is better than A, and A and B are equally good) within the prompt. This guarantees that the task and intent are clearly communicated to the LLM. For specific prompt examples, please see Appendix B.
3.3 Experiment 3: MSMARCO
In the third experiment, a relevance judgment task, we evaluate the models’ ability to identify the most relevant passage for a given query, regardless of the order in which the passages are presented. We use Microsoft’s MSMARCO dataset (Bajaj et al., 2016), [^3] https://huggingface.co/datasets/microsoft/ms_marco which contains queries paired with multiple candidate passages. Each query includes a binary array in which a value of $1$ refers to the index of the most relevant passage and 00 indicates less relevant passages. We shuffle the passages to test whether the models consistently classify the same passage as the most relevant.
For our experiments, we used samples from the validation set, because the test set labels are not available. From the $101093$ entries in the validation set, we first filtered for those in which a most relevant passage was determined (i.e., the binary array had a cell with a value of $1$). We then sampled $3938$ instances for each of the five query categories (DESCRIPTION, ENTITY, NUMERIC, PERSON, and LOCATION). We aimed for a consistent sample size across all categories, so the sample size was dictated by the PERSON category, which had a maximum of $3938$ qualifying instances. For this task, which is slightly more complex than the two previous ones, our few-shot prompts include five examples.
3.4 Experiment 4: MMLU
In the fourth experiment, we evaluate the models’ ability to answer multiple-choice questions while assessing their robustness to changes in the order of answer choices. We use the MMLU dataset (Hendrycks et al., 2020), [^4] https://huggingface.co/datasets/cais/mmlu which covers $57$ diverse subjects, including humanities, STEM, social sciences, and other specialized domains, making it a comprehensive test of LLMs’ general knowledge and reasoning capabilities.
@@ 3.4 Experiment 4: MMLU
3.5 Experiment 5: MedMCQA
In the fifth experiment, we evaluate the models’ performance using the MedMCQA dataset (Pal et al., 2022), [^5] https://huggingface.co/datasets/openlifescienceai/medmcqa a multiple-choice dataset designed to assess medical knowledge. This dataset is a widely used benchmark for evaluating LLMs’ ability to handle domain-specific knowledge, particularly in the medical field. We use the $2816$ examples from the validation set because the test set labels are not publicly available. From the validation set, we only select the questions for which the “correct option” field (cop) was not equal to $-1$, indicating that a correct answer exists. Additionally, we filtered the dataset to only include questions where the choice_type field was set to single, ensuring that each question has exactly one correct answer. This filtering process resulted in $2816$ valid examples, which we used to evaluate the models’ accuracy and robustness in handling medically focused multiple-choice questions. Similar to the previous experiment, the few-shot prompts include five examples.
@@ 4 Experimental Results
Table 1 presents our findings for the MRPC task (3.1). The results indicate that shuffling the order of input leads to a meaningful change in performance for both models. In the zero-shot setup, we observe a 2.78% drop in performance for GPT-4o, while GPT-4o mini shows almost no variation. [^6] Exact results with four decimal points reveal a slight drop. In the few-shot setup, GPT-4o experiences a 1.31% decrease, whereas the mini version shows a 1.51% improvement. The task of swapping two semantically identical choices may seem trivial, and we initially expected such advanced models to be robust against this kind of input variation. Surprisingly, however, the results suggest that current LLMs remain sensitive to small changes in input order. Even more unexpectedly, the mini version demonstrates greater stability compared to its more sophisticated counterpart. Ideally, there should be minimal or no change in performance between the original and shuffled orders, but our findings reveal otherwise. This further highlights the need for detailed investigation into the underlying causes of order sensitivity.
Experiment GPT-4o GPT-4o mini
Precision Recall F1 $\Delta$(%) Precision Recall F1 $\Delta$(%)
Zs-O 0.77 0.71 0.74 -2.78 0.75 0.61 0.67 0.00
Zs-S 0.75 0.69 0.72 0.76 0.60 0.67
Fs-O 0.78 0.77 0.77 -1.31 0.7485 0.5751 0.65 +1.51
Fs-S 0.77 0.76 0.76 0.7546 0.5791 0.66
@@
Table 1: MRPC performance comparison for GPT-4o and GPT-4o mini with percentage change in F1 score after modifying the order. The delta ($\Delta$) is calculated by taking the difference between the F1 score after and the F1 score before shuffling, then dividing the difference by the F1 score after shuffling, e.g. $-2.78\%=\frac{0.72-0.74}{0.72}\times 100$. Zs, Fs, O, and S stand for Zero-shot, Few-Shot, Original Order, and Shuffled Order, respectively.
The results obtained from the WebGPT task (3.2) in Table 2 are even more surprising. Although few-shot learning was expected to alleviate the problem by providing more informative prompts, it failed to prevent significant performance degradation. In the zero-shot setting, the decline is even more severe. This can be attributed to the increased complexity of the WebGPT task compared to MRPC, illustrating how LLMs become increasingly fragile as task complexity grows.
Experiment GPT-4o GPT-4o mini
Precision Recall F1 $\Delta$(%) Precision Recall F1 $\Delta$(%)
ZS-O 0.55 0.52 0.53 -6.00 0.61 0.52 0.56 -8.93
ZS-S 0.50 0.50 0.50 0.51 0.51 0.51
FS-O 0.57 0.51 0.54 -3.84 0.60 0.50 0.55 -14.58
FS-S 0.52 0.51 0.52 0.48 0.48 0.48
Table 2: WebGPT results for GPT-4o and GPT-4o mini with percentage change ($\Delta$) in F1 scores.
The length of inputs (i.e. questions and answers) in the WebGPT task is considerably longer than in MRPC, which we believe contributes to the observed results. Across nearly all experiments, we observed a direct correlation between input length and vulnerability to order changes: the longer the prompt, the greater the performance fluctuations. Despite these observations, the extent of the performance deterioration remains disappointing, especially since the only modification introduced was swapping the order of two semantically identical choices.
One possible explanation for this behavior is the auto-regressive nature of LLMs. While this does not necessarily imply that the WebGPT data was seen during training, it suggests that certain linguistic patterns that the models are used to processing in a certain order are disrupted in the shuffled setting, that leads to performance instability.
Table 3 presents MSMARCO results that closely mirror those of the WebGPT task. The longer inputs (one query and multiple passages) in this task pose a greater challenge for LLMs and results in more significant performance drops. While few-shot prompting is expected to improve performance by providing examples, it fails to do so in this experiment. Instead, the additional examples increase the prompt length, ultimately proving counterproductive. This effect is particularly evident in the GPT-4o mini, which is less capable of handling complex and lengthy inputs.
Experiment GPT-4o GPT-4o mini
Precision Recall F1 $\Delta$(%) Precision Recall F1 $\Delta$(%)
ZS-O 0.49 0.49 0.49 -6.66 0.49 0.49 0.48 -11.62
ZS-S 0.47 0.46 0.45 0.45 0.43 0.43
FS-O 0.48 0.48 0.48 -9.09 0.49 0.47 0.47 -11.90
FS-S 0.46 0.44 0.44 0.46 0.42 0.42
@@
Table 3: MSMARCO relevance judgment results for GPT-4o and GPT-4o mini with percentage change ($\Delta$) in F1 scores.
@@
The key question that arises from these results is: why does changing the input order consistently lead to performance degradation? Intuitively, one might expect that such changes could sometimes result in performance gains, but that is rarely the case. This observation further supports the aforementioned hypothesis that LLMs, due to their auto-regressive nature, are accustomed to processing inputs in a specific order, relying either on expected linguistic patterns or, in cases of data contamination, exact words.
Table 4, reports results from our multiple-choice question-answering task on MMLU, and reveals inconsistencies that are challenging to interpret. Performance fluctuations are more pronounced in GPT-4o mini. Interestingly, few-shot prompting benefits GPT-4o mini but harms the GPT-4o. In the zero-shot setup of GPT-4o, no significant performance change is observed between the original and shuffled inputs, though minor degradation is seen when measured with higher precision. While order variations introduce some differences, they are not substantial. This suggests that for more complex tasks (those involving reasoning or analytical aspects), LLMs may focus more on content, which could potentially make them less sensitive to input order. To explore this further, we conducted a deeper investigation.
Experiment GPT-4o GPT-4o mini
Precision Recall F1 $\Delta$(%) Precision Recall F1 $\Delta$(%)
ZS-O 0.83 0.83 0.83 0.00 0.77 0.76 0.76 -4.10
ZS-S 0.83 0.83 0.83 0.74 0.73 0.73
FS-O 0.84 0.84 0.84 -1.20 0.75 0.74 0.74 +1.33
FS-S 0.83 0.83 0.83 0.75 0.75 0.75
@@
Table 4: MMLU multiple-choice question answering results for GPT-4o and GPT-4o mini with percentage change ($\Delta$) in F1 scores.
We examined F1 scores at the category level within MMLU in the hope of uncovering meaningful patterns. We focused on the zero-shot results of GPT-4o and the few-shot results of GPT-4o mini, as these cases deviated from our earlier findings. Specifically, we analyzed which categories experienced performance drops and which saw improvements after shuffling. Categories like abstract_algebra, conceptual_physics, high_school_mathematics, and machine_learning showed declines, whereas categories, such as philosophy, prehistory, professional_accounting, and world_religions showed improvements. Despite these findings, establishing a consistent pattern for the impact of input order remains challenging. Our observations are limited, and as previously noted, the overall results from MMLU remain inconclusive. However, we are cautiously optimistic that as tasks become more complex and require deeper reasoning, LLMs will move beyond surface-level representations toward a more profound understanding of input content. This shift presents opportunities for further research and refinement of LLMs.
Since our multiple-choice results from MMLU were inconclusive, we conducted an experiment with a more complex multiple-choice dataset, MedMCQA, whose results are shown in Table 5. The task proved particularly challenging for GPT-4o mini, where we observed noticeable performance drops after shuffling the input order. Additionally, including examples in the prompts had little impact, as there was almost no difference between zero-shot and few-shot runs. For GPT-4o, performance remained relatively unchanged, with a slightly negative trend. This raises the possibility, as discussed previously, that LLMs may focus more on content for complex tasks, making them less sensitive to input order variations.
Experiment GPT-4o GPT-4o mini
Precision Recall F1 $\Delta$(%) Precision Recall F1 $\Delta$(%)
ZS-O 0.7696 0.7667 0.7681 -0.85 0.67 0.66 0.66 -3.12
ZS-S 0.7633 0.7601 0.7616 0.65 0.64 0.64
FS-O 0.7606 0.7564 0.7584 -0.17 0.66 0.66 0.66 -3.12
FS-S 0.7593 0.7550 0.7571 0.64 0.64 0.64
@@
Table 5: MedMCQA multiple-choice question answering results for GPT-4o and GPT-4o mini with percentage change ($\Delta$) in F1 scores.
4.1 Summary of Observations
The findings from our experiments (a subset of which were presented in this paper) highlight the following key insights:
  • The longer the input, the harder it becomes for LLMs to process, and these challenging situations increasingly make LLMs vulnerable to performance drops when the input order is altered.

### v2

# The Order Effect: Investigating Prompt Sensitivity to Input Order in LLMs

@@ Abstract
As large language models (LLMs) become integral to diverse applications, ensuring their reliability under varying input conditions is crucial. One key issue affecting this reliability is order sensitivity, wherein slight variations in the input arrangement can lead to inconsistent or biased outputs. Although recent advances have reduced this sensitivity, the problem remains unresolved. This paper investigates the extent of order sensitivity in LLMs whose internal components are hidden from users (such as closed-source models or those accessed via API calls). We conduct experiments across multiple tasks, including paraphrasing, relevance judgment, and multiple-choice questions. Our results show that input order significantly affects performance across tasks, with shuffled inputs leading to measurable declines in output accuracy. Few-shot prompting demonstrates mixed effectiveness and offers partial mitigation; however, fails to fully resolve the problem. These findings highlight persistent risks, particularly in high-stakes applications, and point to the need for more robust LLMs or improved input-handling techniques in future development.
@@ 1 Introduction
One necessary requirement for reasoning is order independence. A model should provide the same consistent response to a query regardless of the order of its content. Historically, LLMs have struggled with this issue. Swapping subsequences within semantically identical inputs often leads to significant changes in output, a problem that worsens as inputs grow in size and complexity (He et al., 2024). Today, improvements in LLMs promise more accurate responses that mitigate order dependency. However, it still remains unclear whether these improvements are sufficient to reduce order issues when these models are used in the wild.
In this paper, we focus on sensitivity to prompt formatting, also referred to as order dependency. This problem has been previously explored in the context of multiple-choice questions (Pezeshkpour & Hruschka, 2023; Zheng et al., 2024), laying the foundation for our research. Building on this foundation, our aim is to provide a fresh perspective and expand the analysis with additional data points and newer models to investigate the problem more thoroughly. Although the issue may seem trivial or some may question the need for further study, we demonstrate that it persists, continues to cause problems, and warrants ongoing investigation. We hope to uncover patterns or root causes that can help mitigate order dependency in the future. To demonstrate the severity of the problem, we conduct a simple experiment in which we submit the following prompt twice to GPT-4o:
@@
To investigate this problem, we experiment with GPT-4o, GPT-4o mini, and DeepSeek (the R1-Distill-Llama-70B version) (Liu et al., 2024) and measure their performance on various order-sensitive tasks. In our setup, we assume that the internal components of these models are hidden from users or that their architectures are not user-modifiable. For GPT models, this assumption holds as they are widely used by many users, yet we have no clear insight into their internal workings. In addition to the GPT models, we include DeepSeek. Although DeepSeek is released as an open-source solution, non-technical users may be misled by this term. In practice, such models are often accessed via APIs or embedded into applications, where users only consume their outputs and have little influence over their responses or internal mechanisms.
As LLMs become increasingly embedded in everyday applications, many users might lack the technical expertise or simply the interest to modify or improve them. This can lead to the unintentional spread of inaccurate information across the internet, public discourse, and even scientific literature. While technically skilled individuals and teams can mitigate such issues (e.g., by modifying open-source code or applying pre-/post-processing filters to LLM inputs/outputs), this paper assumes models are used as-is, by users who may not have the ability or motivation to intervene. Although similar issues have been observed in open-source models, sometimes even more severe, we exclude them from our scope. Open-source models are primarily developed by technical users in more controlled environments, whereas closed-source models are more accessible, making it essential to understand their potential inconsistencies. We hope this work contributes to the broader conversation on the reliability of LLMs as decision-making tools in complex, real-world settings.
@@ 3 Methodology and Experimental Design
To study the impact of order, we designed five experiments, each consisting of four sub-experiments. In the first sub-experiment, we evaluate the models’ performance in a zero-shot setting using the original order of questions and corresponding choices, where the LLM’s task is to select the correct option. In the second sub-experiment, we again assess the models’ zero-shot performance, but this time the entries are presented in a randomized order. For example, a question with the original order of choices, such as “a” and “b”, will be presented in its new form with the choices reversed, namely “b” and “a” (see the following sections for prompt examples).
@@
Finally, in the fourth sub-experiment, we assess the models’ performance in a few-shot setting where the order of entries is randomized. In the second and fourth sub-experiments, reordering does not follow any specific pattern. Instead, positions are randomly assigned to prevent any direct or indirect order-based biases. Together, these four setups allow us to compare model performance in zero-shot and few-shot settings and evaluate the impact of input order on their outputs.
@@ 3.1 Experiment 1: MRPC
Where sentence_1 and sentence_2 are variables which are replaced with the real sentences shared above. This prompt is used for the zero-shot setup with the original order. For the zero-shot setup with the shuffled order, the only change we make to the prompt is swapping the order of the sentences. Specifically, the part of the prompt that needs to be modified is shown below:
@@
3.2 Experiment 2: MSMARCO
In the second experiment, a relevance judgment task, we evaluate the models’ ability to identify the most relevant passage for a given query, regardless of the order in which the passages are presented. We use Microsoft’s MSMARCO dataset (Bajaj et al., 2016), [^2] https://huggingface.co/datasets/microsoft/ms_marco which contains queries paired with multiple candidate passages. Each query includes a binary array in which a value of $1$ refers to the index of the most relevant passage and 00 indicates less relevant passages. We shuffle the passages to test whether the models consistently classify the same passage as the most relevant.
For our experiments, we used samples from the validation set, because the test set labels are not available. From the $101093$ entries in the validation set, we first filtered for those in which a most relevant passage was determined (i.e., the binary array had a cell with a value of $1$). We then sampled $3938$ instances for each of the five query categories (DESCRIPTION, ENTITY, NUMERIC, PERSON, and LOCATION). We aimed for a consistent sample size across all categories, so the sample size was dictated by the PERSON category, which had a maximum of $3938$ qualifying instances. For this task, which is slightly more complex than the previous one, our few-shot prompts include five examples.
3.3 Experiment 3: MMLU
In the third experiment, we evaluate the models’ ability to answer multiple-choice questions while assessing their robustness to changes in the order of answer choices. We use the MMLU dataset (Hendrycks et al., 2020), [^3] https://huggingface.co/datasets/cais/mmlu which covers $57$ diverse subjects, including humanities, STEM, social sciences, and other specialized domains, making it a comprehensive test of LLMs’ general knowledge and reasoning capabilities.
@@ 3.3 Experiment 3: MMLU
3.4 Experiment 4: MedMCQA
In the fourth experiment, we evaluate the models’ performance using the MedMCQA dataset (Pal et al., 2022), [^4] https://huggingface.co/datasets/openlifescienceai/medmcqa a multiple-choice dataset designed to assess medical knowledge. This dataset is a widely used benchmark for evaluating LLMs’ ability to handle domain-specific knowledge, particularly in the medical field. We use the $2816$ examples from the validation set because the test set labels are not publicly available. From the validation set, we only select the questions for which the “correct option” field (cop) was not equal to $-1$, indicating that a correct answer exists. Additionally, we filtered the dataset to only include questions where the choice_type field was set to single, ensuring that each question has exactly one correct answer. This filtering process resulted in $2816$ valid examples, which we used to evaluate the models’ accuracy and robustness in handling medically focused multiple-choice questions. Similar to the previous experiment, the few-shot prompts include five examples.
3.5 Experiment 5: WebGPT
The fifth and last experiment focuses on comparison consistency. We evaluate the models’ ability to compare two answers to a given question and determine which answer a human judge would prefer, regardless of the order in which the answers are presented. We use OpenAI’s WebGPT dataset (Nakano et al., 2021), [^5] https://huggingface.co/datasets/openai/webgpt_comparisons which provides pairwise comparisons derived from a reward model trained on human feedback to reflect real-world preferences for long-form question answering. To capture the models’ behavior, we switch the order of the answers and observe whether they consistently select the same preferred answer. We use the $1958$ examples from the test set. Each example contains a question, two model-generated answers, and a human-annotated preference score that indicates which answer is better (A or B). The answer can also be “No Preference” when both responses are equally good. To ensure that the prompt in the few-shot setting is informative, we include one example from each case (A is better than B, B is better than A, and A and B are equally good) within the prompt. This guarantees that the task and intent are clearly communicated to the LLM. For specific prompt examples, please see Appendix B.
@@ 4 Experimental Results
Table 1 presents our findings for the MRPC task (3.1). The results indicate that shuffling the order of input leads to a performance drop for all models. Even for cases whose delta is zero, unrounded percentage shows a slight decline. In the zero-shot setup, we observe a 2.77% drop in for GPT-4o. GPT-4o mini and DeepSeek show a similar trend. In the few-shot setup, the gap is smaller that indicates extra context might be useful but the overall trend stays unchanged. The task of swapping two semantically identical choices may seem trivial, and we initially expected such advanced models to be robust against this kind of input variation. Surprisingly, however, the results suggest that current LLMs remain sensitive to small changes in the input order. Even more unexpectedly, the mini version demonstrates greater stability compared to a more sophisticated counterpart. Ideally, there should be minimal or no change in performance between the original and shuffled orders, but our findings reveal otherwise. The performance further declines on other datasets, which highlights the need for a detailed investigation into the underlying causes of order sensitivity.
Setup GPT-4o GPT-4o mini DeepSeek
P R F1 $\Delta$ P R F1 $\Delta$ P R F1 $\Delta$
Zs-O 0.77 0.71 0.72 -2.77 0.75 0.61 0.61 -1.63 0.79 0.73 0.73 -1.36
Zs-S 0.75 0.69 0.70 0.76 0.60 0.60 0.77 0.72 0.72
Fs-O 0.78 0.77 0.77 -1.29 0.75 0.58 0.57 0.0 0.76 0.69 0.70 0.0
Fs-S 0.77 0.76 0.76 0.75 0.58 0.57 0.77 0.69 0.70
@@
Table 1: MRPC performance comparison for GPT-4o, GPT-4o mini, and DeepSeek with percentage change in F1 score after modifying the order. The delta ($\Delta$) is calculated by taking the difference between the F1 score after and the F1 score before shuffling, then dividing the difference by the F1 score after shuffling, e.g. $-2.77\%=\frac{0.70-0.72}{0.72}\times 100$. Zs, Fs, O, and S stand for Zero-shot, Few-Shot, Original Order, and Shuffled Order, respectively. P and R are acronyms for precision and recall and F1 is a class-based, weighted average.
Table 2 presents the results on the MSMARCO dataset. We considered the binary decision-making scenario in MRPC to be too trivial for studying the problem in depth. Therefore, we aimed to increase the task complexity and input length to observe how models behave in a more challenging setting. The MSMARCO dataset (involves one query paired with multiple passages), obviously, presents longer inputs that poses a greater challenge for LLMs and this leads to sever declines in quality of responses. Few-shot prompting was also expected to enhance performance by providing illustrative examples, however, it fails to do so in this case and even worsens the outcome. This may be due to the increased prompt length introduced by the examples, which ultimately becomes counterproductive. Additionally, the weaker models exhibit greater sensitivity compared to the GPT models.
Setup GPT-4o GPT-4o mini DeepSeek
P R F1 $\Delta$ P R F1 $\Delta$ P R F1 $\Delta$
Zs-O 0.49 0.49 0.49 -6.12 0.50 0.49 0.49 -12.24 0.43 0.45 0.43 -2.32
Zs-S 0.47 0.46 0.46 0.46 0.43 0.43 0.43 0.44 0.42
Fs-O 0.49 0.48 0.48 -8.33 0.50 0.48 0.47 -10.63 0.43 0.46 0.43 -13.95
Fs-S 0.46 0.45 0.44 0.46 0.43 0.42 0.39 0.39 0.37
@@
Table 2: MSMARCO relevance judgment results for GPT-4o, GPT-4o mini, and DeepSeek with percentage change ($\Delta$) in F1 scores.
@@
So far we have investigated two datasets with different behavior, and the key question that arises from the results is: why does changing the input order consistently lead to performance degradation? Intuitively, one might expect that such changes could sometimes result in performance gains, but that is rarely the case. This observation might hint at a hypothesis that LLMs, due to their auto-regressive nature, are accustomed to processing inputs in a specific order, relying either on expected linguistic patterns or, in cases of data contamination, exact words.
Setup GPT-4o GPT-4o mini DeepSeek
P R F1 $\Delta$ P R F1 $\Delta$ P R F1 $\Delta$
Zs-O 0.84 0.83 0.83 0.0 0.77 0.76 0.76 -2.63 0.78 0.65 0.66 -6.06
Zs-S 0.83 0.83 0.83 0.74 0.74 0.74 0.78 0.61 0.62
Fs-O 0.85 0.85 0.85 -1.17 0.76 0.75 0.75 0.0 0.80 0.60 0.67 -4.47
Fs-S 0.84 0.84 0.84 0.76 0.75 0.75 0.81 0.62 0.64
@@
Table 3: MMLU multiple-choice question answering results for GPT-4o, GPT-4o mini, and DeepSeek with percentage change ($\Delta$) in F1 scores.
Table 3 presents the results from our multiple-choice question-answering task on the MMLU dataset, which features a completely different type of input. Performance fluctuations are more pronounced in DeepSeek, although performance degradation is also observed in the GPT models. Interestingly, few-shot prompting benefits the smaller model, GPT-4o mini, but negatively affects GPT-4o. This may be because adding extra examples in few-shot prompting strengthens the reasoning ability of smaller models, while the same approach could introduce irrelevant or distracting information that misleads more powerful models, which are already capable of solving the task without additional guidance. Nonetheless, the GPT models demonstrate relatively high performance in both zero-shot and few-shot settings.
We also examined F1 scores at the category level within the MMLU dataset, aiming to uncover meaningful patterns. Our analysis focused on GPT-4o and the few-shot results of GPT-4o mini. DeepSeek exhibited performance degradation across all settings that makes it less informative for this comparison, but, the selected configurations for the GPT models diverge from our earlier observations, that drew our attention for further investigation. We analyze which categories experienced performance drops and which saw improvements after input shuffling. Notably, categories such as abstract_algebra, conceptual_physics, high_school_mathematics, and machine_learning showed performance declines, whereas others such as philosophy, prehistory, and world_religions showed improvements.
Despite these observations, identifying a consistent pattern in how input order affects performance remains difficult. However, it appears that for text-based categories (e.g. philosophy) involving reading comprehension, LLMs are more sensitive to input order. In contrast, for complex, reasoning-intensive tasks (e.g. algebra), LLMs may be more resilient by potentially moving beyond surface-level representations and toward a deeper understanding of the input content.
Since our multiple-choice results on the MMLU dataset were not quite conclusive, we conducted an additional experiment using a more complex multiple-choice dataset, MedMCQA, in the hope of gaining further insights. The results of this experiment are presented in Table 4. The findings are largely consistent with those from the MMLU task: GPT-4o significantly outperforms both its smaller variant, GPT-4o mini, and DeepSeek, with the latter showing the weakest performance overall.
Setup GPT-4o GPT-4o mini DeepSeek
P R F1 $\Delta$ P R F1 $\Delta$ P R F1 $\Delta$
Zs-O 0.77 0.77 0.77 -1.29 0.67 0.67 0.67 -2.98 0.69 0.64 0.64 -6.25
Zs-S 0.76 0.76 0.76 0.66 0.65 0.65 0.67 0.59 0.60
Fs-O 0.76 0.76 0.76 0.0 0.66 0.66 0.66 -3.03 0.65 0.62 0.62 -8.06
Fs-S 0.76 0.76 0.76 0.65 0.64 0.64 0.65 0.57 0.57
@@
Table 4: MedMCQA multiple-choice question answering results for GPT-4o, GPT-4o mini, and DeepSeek with percentage change ($\Delta$) in F1 scores.
All the experiments reported so far consistently demonstrate that changing the input order generally leads to performance degradation. This raised an important question for us: can altering the order ever improve performance? We looked in multiple datasets and WebGPT provided such a case. Results obtained from the WebGPT task (3.5) are reported in Table 5. Except for the few-shot setting of DeepSeek, all settings showed consistent performance improvements. One possible explanation could be that WebGPT is in fact specifically designed to fine-tune and improve LLMs for these types of issues. For the same reason, there is also a possibility that LLMs were exposed to this dataset during training. However, setting aside such speculation, we analyzed the results to understand where and why improvements occurred but we did not observe any consistent patterns.
Setup GPT-4o GPT-4o mini DeepSeek
P R F1 $\Delta$ P R F1 $\Delta$ P R F1 $\Delta$
Zs-O 0.55 0.52 0.48 2.08 0.61 0.52 0.46 10.86 0.13 0.36 0.19 21.05
Zs-S 0.50 0.50 0.49 0.51 0.51 0.51 0.58 0.36 0.23
Fs-O 0.57 0.51 0.46 10.86 0.60 0.50 0.45 6.66 0.13 0.35 0.19 -5.26
Fs-S 0.52 0.51 0.51 0.48 0.48 0.48 0.55 0.33 0.18
Table 5: WebGPT results for GPT-4o, GPT-4o mini, and DeepSeek with percentage change ($\Delta$) in F1 scores.
4.1 Summary of Findings
While no clear set of patterns emerged to fully explain the results, a few weak trends were observed:
  • The longer the input, the more difficult it becomes for LLMs to process effectively, and such complex inputs appear to increase the models’ vulnerability to performance degradation when the input order is altered.